1. A fair coin is flipped 6 times. (a) Find the probability that there is exactly...

1.

A fair coin is flipped

6

times.

(a)

Find the probability that there is exactly

3

heads.

(b)

Find the probability the there are at least

3

heads.

(c)

Find the probability that there is no head.

(d)

Find the expected number of heads.

(e)

Find the standard deviation of the number of heads.
Answer #1

Coin flipped 6 times so total outcomes =2^6 = 32

b) P(at least 3) = P(3) + P(4) + P(5) + P(6)

= 21/32 or 0.65625

d) Mean = n*p = (6)*(1/2) = 3